Grace is twice as old as Janice. seven years hence, the sum of their ages will be 44 years. How old are they?
Answer by JulietG(1812)

G = 2J
(G+7)+(J+7)=44
Substitute the known value of G from the first equation into the second
(2J+7)+(J+7)=44
3J+14 = 44
Subtract 14 from each side
3J = 30
Divide each side by 3
J = 10


If Janice is 10, then Grace is 20.
In 7 years, Janice will be 17 and Grace will be 27.
17 + 27 = 44